OpenStudy (anonymous):

Estimate 48% of $23.95

OpenStudy (anonymous):

11.50

OpenStudy (anonymous):

rounded

OpenStudy (anonymous):

okay thank you :}

OpenStudy (anonymous):

do u know how to get that? in case another problem like that comes up again?

OpenStudy (alexandervonhumboldt2):

23.95*48/100=solve

OpenStudy (misty1212):

i would round 48 % to one half, and 23.95 to 24 and say half of twenty four is 12

OpenStudy (misty1212):

it does say "estimate" right? not "solve using a calculator" or sommat

OpenStudy (photon336):

@emma_music13 \[\frac{ x }{ 23.95 } = \frac{ 48 }{ 100 }\]

OpenStudy (mathmale):

Photon: Makes great sense, EXCEPT that the instructions ask for an ESTIMATE.